OpenStudy (anonymous):

An arithmetic sequence is defined by the recursive formula t1 = -30, tn + 1 = tn - 4, where n ∈N and n ≥ 1. Which is the general term of the sequence? A) tn = -30 - 4n, where n ∈N and n ≥ 1 B) tn = -30 + 4n, where n ∈N and n ≥ 1 C) tn = -26 - 4n, where n ∈N and n ≥ 1 D) tn = -26 + 4n, where n ∈N and n ≥ 1

OpenStudy (amoodarya):

t2=t1-4 so t2=-30-4=-34 so d=-4 t(n)=t1 +(n-1)d t(n)=-30 +(n-1)(-4) =-30-4n +4 so t(n)=-26-4n for all n>=1